If velocity ,time and force were chosen as basic quanties ,find the dimensions of mass
1
Expert's answer
2020-01-27T11:10:54-0500

Solution. From Newton’s second law


"F=ma"

where F is force; m is mass; a is acceleration.

By definition of acceleration


"a=\frac {\Delta v}{\Delta t}"

where Δv is speed change; Δt is time. Substituting the second formula in the first we get


"F=m\frac {\Delta v}{\Delta t} \implies m= \frac {F \Delta t} {\Delta v}=F \Delta t(\Delta v)^{-1}"

Thus, the dimension of mass is


"[m]=N \cdot s^2 \cdot m^{-1}"